OpenStudy (anonymous):

simplify each expression. (4m)^5

OpenStudy (stamp):

Didn't somebody just explain these to you?

OpenStudy (mathstudent55):

To raise a product to a power, raise each factor to the power. Then multiply the factors together.

OpenStudy (stamp):

If not, review the blue table. http://tutorial.math.lamar.edu/Classes/Alg/IntegerExponents.aspx Property 4.

OpenStudy (anonymous):

The answer is 1024m^5

OpenStudy (mathstudent55):

For example: (3x)^4 = 3^4 * x^4 = 81x^4 Now apply this to your problem.